# Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?

### v4

@@ Appendix E Technical Details
The backtesting framework and traditional rule-based strategies in FINSABER are implemented using BackTrader [^9] https://www.backtrader.com/ and Papers With Backtest [^10] https://paperswithbacktest.com/. Reinforcement learning-based methods are implemented using FinRL (liu2021finrl). FINSABER supports two operational modes: “LLM” mode and “BT” mode. The “LLM” mode is tailored for strategies that leverage multi-modal inputs, including financial news and regulatory filings. In contrast, the “BT” mode is built directly on BackTrader, offering robust support for traditional rule-based strategies while maintaining a familiar interface to facilitate easy migration from standard BackTrader workflows.
@@ Appendix F Comparative Drawdown Analysis via Underwater Plots
The plots are derived by calculating the percentage loss of a portfolio’s equity curve from its running maximum value (its previous peak). At any given point in time, the drawdown $D_{t}$ is calculated as:
(8) $D_{t}=\frac{\text{Current Value}_{t}-\text{Previous Peak}_{t}}{\text{Previous Peak}_{t}}$
@@
The third row, depicting the 2008 Global Financial Crisis, provides the most critical insight into the agents’ flaws. While a single stock is expected to be more volatile than the index during a crash, the LLM strategies, particularly FinMem, catastrophically amplify this downside risk. For DE, the FinMem strategy’s drawdown approaches -75%, a far more severe loss than the SPX benchmark’s -50%. Rather than providing any form of risk mitigation, the agents appear to make pro-cyclical decisions that accelerate losses. The FinAgent strategy, true to its more conservative nature, often mitigates some of these losses relative to FinMem, yet it still fails to generate a positive outcome. For instance, while its drawdown for CNX is shallower than FinMem’s, it remains severe and prolonged. This relative outperformance is insufficient and aligns with our market regime analysis (§7), which finds that both agents are poorly calibrated for bear markets and ultimately succumb to losses (cont2001empirical).

### v5

@@ Appendix E Technical Details
The backtesting framework and traditional rule-based strategies in FINSABER are implemented using BackTrader [^9] https://www.backtrader.com/ and Papers With Backtest [^10] https://paperswithbacktest.com/. Reinforcement learning-based methods are implemented using FinRL (Liu et al., 2021). FINSABER supports two operational modes: “LLM” mode and “BT” mode. The “LLM” mode is tailored for strategies that leverage multi-modal inputs, including financial news and regulatory filings. In contrast, the “BT” mode is built directly on BackTrader, offering robust support for traditional rule-based strategies while maintaining a familiar interface to facilitate easy migration from standard BackTrader workflows.
@@ Appendix F Comparative Drawdown Analysis via Underwater Plots
The plots are derived by calculating the percentage loss of a portfolio’s equity curve from its running maximum value (its previous peak). At any given point in time, the drawdown $D_{t}$ is calculated as: $D_{t}=(\text{Current Value}_{t}-\text{Previous Peak}_{t})/\text{Previous Peak}_{t}$.
@@
The third row, depicting the 2008 Global Financial Crisis, provides the most critical insight into the agents’ flaws. While a single stock is expected to be more volatile than the index during a crash, the LLM strategies, particularly FinMem, catastrophically amplify this downside risk. For DE, the FinMem strategy’s drawdown approaches -75%, a far more severe loss than the SPX benchmark’s -50%. Rather than providing any form of risk mitigation, the agents appear to make pro-cyclical decisions that accelerate losses. The FinAgent strategy, true to its more conservative nature, often mitigates some of these losses relative to FinMem, yet it still fails to generate a positive outcome. For instance, while its drawdown for CNX is shallower than FinMem’s, it remains severe and prolonged. This relative outperformance is insufficient and aligns with our market regime analysis (§7), which finds that both agents are poorly calibrated for bear markets and ultimately succumb to losses (Cont, 2001).